# Variational Supervised Contrastive Learning

### v2

@@ 4.2 Few-Shot Learning Performance
Given the practical importance of learning from limited labeled data, we evaluate VarCon in few-shot learning scenarios by comparing it with SupCon across different per-class sample sizes. Table 2 presents this comparison on ImageNet using ResNet-50, with sample sizes ranging from extremely limited ($N=50$) to moderate ($N=1000$). Results are averaged across five random subsets per configuration to account for sampling variability. VarCon consistently outperforms SupCon across all sample sizes, with advantages more pronounced in data-scarce settings: With 100 samples per class, VarCon achieves 37.81% Top-1 accuracy, surpassing SupCon (36.57%) by 1.24%. This advantage persists at $N=200$ (51.10% vs. 50.25%) and $N=500$ (65.83% vs. 64.91%). As training data increases, VarCon maintains superior performance, though the gap narrows slightly (73.21% vs. 73.04% at $N=1000$). These results suggest our variational framework provides greatest benefit when training data is scarce—where robust representation learning is most challenging. By modeling feature uncertainty explicitly, VarCon learns more effective representations from limited data, making it well-suited for real-world applications where large labeled datasets are often unavailable.
4.3 Transfer Learning Performance
@@
Method Food CIFAR10 CIFAR100 Birdsnap SUN397 Cars Aircraft VOC2007 DTD Pets Caltech-101 Flowers Mean
SimCLR-50 88.21 97.62 85.86 75.91 63.52 91.33 87.40 83.98 73.23 89.22 92.11 97.02 85.45
SupCon-50 87.28 97.43 84.26 75.20 58.03 91.71 84.08 85.18 74.82 93.45 91.07 96.08 84.88
VarCon-50 87.31 97.65 84.39 75.27 57.96 91.79 84.25 85.07 74.79 93.52 91.19 96.14 84.94
SupCon-200 88.65 98.32 87.25 76.27 60.44 91.83 88.53 85.12 74.62 93.09 94.87 96.94 86.33
VarCon-200 88.68 98.34 87.30 76.29 60.39 91.85 88.61 85.02 74.47 93.26 94.95 96.96 86.34
4.4 Effect of Data Augmentation and Encoder Architecture
@@ 4.5 Ablation Studies
KNN Classification Analysis. To directly evaluate embedding quality without additional parameterized classifiers, we employed K-nearest neighbor classification, which better reflects the intrinsic structure of the feature space. Figure 3(a) shows that VarCon consistently outperforms SupCon across all K values on ImageNet. For Top-1 accuracy, VarCon’s advantage increases with larger K values (from 77.4% vs. 77.32% at K=5 to 78.07% vs. 77.86% at $K=80$). For Top-5 accuracy, VarCon maintains a consistent 0.27% advantage (83.88% vs. 83.61%) across all K values. This superiority confirms that VarCon produces embedding spaces with clearer decision boundaries and better structured class relationships.
@@ 4.6 Robustness to Image Corruption
To evaluate robustness to real-world image degradation, we tested VarCon on ImageNet-C, which applies 15 different corruption types at 5 increasing severity levels. Figure 3(c) shows Top-1 accuracy on clean ImageNet (“test”) and across all corruption severity levels. VarCon consistently outperforms SupCon with both ResNet-50 and ResNet-200 architectures. With ResNet-50, VarCon maintains a performance advantage ranging from 0.3% to 1.2% as corruption severity increases from level 1 to 5. This graceful degradation pattern can be attributed to our probabilistic framework, which explicitly models feature uncertainty. By representing features as distributions rather than points, VarCon better accommodates input variations caused by corruptions. These results highlight that our variational approach not only performs well on clean data but also demonstrates enhanced robustness to low-quality inputs—a valuable property for real-world applications.
@@ 5 Conclusion
In this work, we introduced Variational Supervised Contrastive Learning (VarCon), a probabilistically grounded framework that reformulates supervised contrastive learning through variational inference over latent class variables. By deriving and maximizing a posterior-weighted evidence lower bound, VarCon overcomes key limitations of conventional contrastive approaches: it explicitly regulates embedding distributions through a principled KL divergence term, replaces exhaustive pairwise comparisons with efficient class-level interactions, and employs confidence-adaptive temperature scaling to precisely control intra-class dispersion. Our extensive empirical evaluation demonstrates that VarCon consistently outperforms leading contrastive methods across diverse benchmarks, achieving superior classification accuracy on CIFAR-10/100 and ImageNet while requiring fewer epochs for convergence. Moreover, VarCon exhibits enhanced robustness to hyperparameter choices, reduced dependency on large batch sizes, stronger performance in low-data regimes, and improved resilience against image corruptions—all while maintaining better transferability across domains. Beyond the empirical advantages, VarCon bridges the theoretical gap between distinguishing and generative paradigms by endowing contrastive objectives with explicit likelihood semantics.
@@ Appendix A Discussion
A.1 Broader Impact
@@ B.2 Extended Evaluations
Self-supervised SimCLR $75.30_{\pm 0.05}$
@@
Supervised Cross-Entropy $83.47_{\pm 0.06}$
@@
On this more rigorous dataset, VarCon achieves 84.12% Top-1 accuracy, outperforming all baseline methods by a statistically significant margin. The performance gap between VarCon and the strongest self-supervised method (SwAV at 81.56%) is significant, showing a solid 2.56% improvement, suggesting that our probabilistic formulation effectively leverages label information to learn more discriminative representations that better align with human perception of image content.
@@ B.3 Embedding Space Evolution Analysis
The visualization demonstrates clear progression from loosely organized clusters at epoch 50 (KNN classifier: 52.74%) to highly structured, well-separated semantic groups by epoch 200 (79.11%). The most significant improvement occurs between epochs 100-150, where the KNN classifier performance jumps from 59.57% to 71.12%, corresponding to the convergence of our ELBO-derived loss function. During this phase, the KL divergence term effectively aligns the auxiliary posterior with the model’s class posterior while ensuring correct class assignments. By epoch 200, VarCon achieves optimal embedding organization with distinct, compact clusters and clear decision boundaries. We employ KNN classification to directly evaluate the quality of pretrained representations without introducing additional trainable parameters that could mask the intrinsic discriminative power of the learned embeddings. The high KNN classifier performance (79.11%) demonstrates that our confidence-adaptive temperature mechanism successfully provides fine-grained control over intra-class dispersion, creating embedding spaces where semantic similarity corresponds to geometric proximity and enabling effective classification through nearest-neighbor search.
@@
Figure 4: Progressive evolution of VarCon embedding space visualization through t-SNE [9] during training on ImageNet validation set. Our variational formulation demonstrates systematic improvement in semantic organization, with KNN-classifier accuracy increasing from 52.74% at epoch 50 to 79.11% at epoch 200 as clusters become increasingly well-separated and semantically coherent. The confidence-adaptive temperature mechanism enables fine-grained control over intra-class dispersion, resulting in embedding spaces with clear decision boundaries and hierarchical semantic structure that facilitate effective nearest-neighbor classification without additional parameterized classifiers.
@@ B.4 Hierarchical Clustering Results
Table 6 presents comprehensive clustering quality metrics. VarCon consistently outperforms SupCon across all evaluation criteria. The Adjusted Rand Index (ARI), which measures the similarity between cluster assignments and ground-truth labels while correcting for chance, increases from 0.613 to 0.634. This improvement suggests that the discovered clusters align more closely with true semantic categories. Similarly, the purity metric, which quantifies the fraction of samples in each cluster belonging to the most common class, improves from 0.755 to 0.774, indicating that individual clusters contain fewer mixed-class samples. The Normalized Mutual Information (NMI) score, measuring the mutual dependence between cluster assignments and true labels, increases from 0.888 to 0.895. This gain reflects a stronger statistical relationship between the discovered structure and semantic categories. The V-measure, which provides the harmonic mean of homogeneity and completeness, shows balanced improvements (0.895 vs. 0.888). Specifically, homogeneity, measuring whether each cluster contains only members of a single class, improves from 0.886 to 0.893, while completeness, measuring whether all members of a class are assigned to the same cluster, increases from 0.890 to 0.896. These metric improvements align with our theoretical framework. The confidence-adaptive temperature mechanism allows VarCon to apply stronger supervision on difficult samples while relaxing constraints on well-classified examples. This adaptive behavior, combined with the KL divergence term that prevents distributional collapse, appears to preserve more nuanced semantic relationships during training. The clustering analysis thus provides empirical support for our variational formulation’s ability to maintain meaningful structure in the representation space.
@@ Appendix C Theoretical Analysis
[…]
@@
×$\quad\times\left\{\sum_{r^{\prime}\neq r}\log\frac{q_{\phi}(r^{\prime}\mid\bm{z})}{p_{\theta}(r^{\prime}\mid\bm{z})}-(C-1)\log\frac{q_{\phi}(r\mid\bm{z})}{p_{\theta}(r\mid\bm{z})}\right\}.$ (46)
@@
  • For confident samples ($p_{\theta}(r\mid\bm{z})>0.5$): The factor $[2p_{\theta}(r\mid\bm{z})-1]>0$. When $p_{\theta}(r\mid\bm{z})$ is high (confident prediction), $\tau_{2}$ approaches $\tau_{1}+\epsilon$, making $q_{\phi}$ more uniform. The gradient direction is determined by the aggregate log-ratio term $\sum_{r^{\prime}\neq r}\log\frac{q_{\phi}(r^{\prime}\mid\bm{z})}{p_{\theta}(r^{\prime}\mid\bm{z})}-(C-1)\log\frac{q_{\phi}(r\mid\bm{z})}{p_{\theta}(r\mid\bm{z})}$. This uniform $q_{\phi}$ typically satisfies $q_{\phi}(r\mid\bm{z})<p_{\theta}(r\mid\bm{z})$ for the ground-truth class and $q_{\phi}(r^{\prime}\mid\bm{z})>p_{\theta}(r^{\prime}\mid\bm{z})$ for other classes, making the aggregate term positive. Consequently, the gradient pushes $\epsilon$ to increase, which in turn increases $\tau_{2}$, making $q_{\phi}$ even more uniform and effectively reducing supervision strength for already well-classified samples. This self-reinforcing mechanism prevents overfitting on easy samples while maintaining stable learning dynamics.
  • For difficult samples ($p_{\theta}(r\mid\bm{z})<0.5$): The factor $[2p_{\theta}(r\mid\bm{z})-1]<0$. When the model is uncertain, $\tau_{2}$ approaches $\tau_{1}-\epsilon$, creating a sharper distribution. The gradient behavior reverses, causing $\epsilon$ to adjust such that $\tau_{2}$ decreases. This reduction in $\epsilon$ leads to a smaller $\tau_{2}$, creating a sharper $q_{\phi}$ distribution that provides stronger supervision precisely where the model needs it most. The same aggregate log-ratio term that increases $\epsilon$ for confident samples now decreases it due to the negative sign factor. The magnitude of this effect scales with the degree of distributional mismatch, ensuring proportional adaptation to the model’s uncertainty.

### v3

@@ 4.2 Few-Shot Learning Performance
Given the practical importance of learning from limited labeled data, we evaluate VarCon in few-shot learning scenarios by comparing it with SupCon across different per-class sample sizes. Table 2 presents this comparison on ImageNet using ResNet-50, with sample sizes ranging from extremely limited ($N=50$) to moderate ($N=1000$). Results are averaged across five random subsets per configuration to account for sampling variability. VarCon consistently outperforms SupCon across all sample sizes, with advantages more pronounced in data-scarce settings: With 100 samples per class, VarCon achieves 37.81% Top-1 accuracy, surpassing SupCon (36.57%) by 1.24%. This advantage persists at $N=200$ (51.10% vs. 50.25%) and $N=500$ (65.83% vs. 64.91%). As training data increases, VarCon maintains superior performance, though the gap narrows slightly (73.21% vs. 73.04% at $N=1000$). These results suggest our variational framework provides greatest benefit when training data is scarce, where robust representation learning is most challenging. By modeling feature uncertainty explicitly, VarCon learns more effective representations from limited data, making it well-suited for real-world applications where large labeled datasets are often unavailable.
4.3 Effect of Data Augmentation and Encoder Architecture
@@
Method Food CIFAR10 CIFAR100 Birdsnap SUN397 Cars Aircraft VOC2007 DTD Pets Caltech‐101 Flowers Mean
SimCLR‐50 88.21 97.62 85.86 75.91 63.52 91.33 87.40 83.98 73.23 89.22 92.11 97.02 85.45
SupCon‐50 87.28 97.43 84.26 75.20 58.03 91.71 84.08 85.18 74.82 93.45 91.07 96.08 84.88
VarCon‐50 87.31 97.65 84.39 75.27 57.96 91.79 84.25 85.07 74.79 93.52 91.19 96.14 84.94
SupCon‐200 88.65 98.32 87.25 76.27 60.44 91.83 88.53 85.12 74.62 93.09 94.87 96.94 86.33
VarCon‐200 88.68 98.34 87.30 76.29 60.39 91.85 88.61 85.02 74.47 93.26 94.95 96.96 86.34
4.4 Transfer Learning Performance
@@ 4.5 Ablation Studies
Analysis of Embedding Space Organization. To directly evaluate embedding quality without additional parameterized classifiers, we employed K-nearest neighbor classification, which better reflects the intrinsic structure of the feature space. Figure 3(a) shows that VarCon consistently outperforms SupCon across all K values on ImageNet. For Top-1 accuracy, VarCon’s advantage increases with larger K values (from 77.4% vs. 77.32% at K=5 to 78.07% vs. 77.86% at $K=80$). For Top-5 accuracy, VarCon maintains a consistent 0.27% advantage (83.88% vs. 83.61%) across all K values. This superiority confirms that VarCon produces embedding spaces with clearer decision boundaries and better structured class relationships. This improved embedding organization is further confirmed by t-SNE visualizations of the embedding space evolution in Appendix B.3 and by quantitative hierarchical clustering metrics in Appendix B.4.
@@ 4.6 Robustness to Image Corruption
To evaluate robustness to real-world image degradation, we tested VarCon on ImageNet-C, which applies 15 different corruption types at 5 increasing severity levels. Figure 3(c) shows Top-1 accuracy on clean ImageNet (“test”) and across all corruption severity levels. VarCon consistently outperforms SupCon with both ResNet-50 and ResNet-200 architectures. With ResNet-50, VarCon maintains a performance advantage ranging from 0.3% to 1.2% as corruption severity increases from level 1 to 5. This graceful degradation pattern can be attributed to our probabilistic framework, which explicitly models feature uncertainty. By representing features as distributions rather than points, VarCon better accommodates input variations caused by corruptions. These results highlight that our variational approach not only performs well on clean data but also demonstrates enhanced robustness to low-quality inputs, which is a valuable property for real-world applications.
4.7 Adaptive Temperature Analysis
(a) Mean $\tau_{2}$ evolution
(b) $\tau_{2}$ density distributions
Figure 4: Evolution of adaptive temperature $\tau_{2}$ during a full 50-epoch ImageNet training with ResNet-50 ($\epsilon=0.02$, $\tau_{1}=0.1$, batch size 4096). (a) Mean $\tau_{2}$ increases from 0.09378 to 0.10656 over the complete training, indicating systematic confidence growth. (b) Density distributions (epochs 10-50) show rightward shift with initial broadening (std: 0.00961 $\rightarrow$ 0.01092) then stabilization, reflecting heterogeneous confidence development as the model distinguishes easy from hard samples.
To investigate our confidence-adaptive temperature mechanism, we analyze $\tau_{2}$ distribution across ImageNet samples during ResNet-50 training (batch size 4096, $\tau_{2}\in[0.08,0.12]$). We fix $\epsilon=0.02$ and $\tau_{1}=0.1$ to ensure comparable temperature spans across epochs, though $\epsilon$ is learnable in our best configuration. Figure 4 shows systematic evolution: initially (epoch 10), over 25% of samples receive near-maximal gradients with $\tau_{2}$ near lower bound, indicating widespread uncertainty. Mean $\tau_{2}$ increases 13.6% during training while maintaining persistent challenging samples near minimum. This heterogeneous evolution demonstrates differential treatment, strong gradients for difficult samples and relaxed constraints for well-learned ones, emerging naturally from KL divergence and confidence-based temperature interaction. This demonstrates our self-adaptive temperature simultaneously adjusts both target sharpness and gradient magnitude based on per-sample confidence (see Appendix D.4 for comparisons with other regularization methods [szegedy2016rethinking, zhang2017mixup, ghosh2022adafocal, tao2023dual, bahri2021locally, jin2024learning]). Detailed $\tau_{2}$ evolution statistics for fixed $\epsilon$ (Table 7) and learnable $\epsilon$ (Table 8) are in Appendix D.2.
@@ 5 Conclusion
In this work, we introduced Variational Supervised Contrastive Learning (VarCon), a probabilistically grounded framework that reformulates supervised contrastive learning through variational inference over latent class variables. By deriving and maximizing a posterior-weighted evidence lower bound, VarCon overcomes key limitations of conventional contrastive approaches: it explicitly regulates embedding distributions through principled KL divergence, replaces exhaustive pairwise comparisons with efficient class-level interactions, and employs confidence-adaptive temperature scaling to control intra-class dispersion. Our extensive evaluation demonstrates VarCon consistently outperforms leading contrastive methods across diverse benchmarks, achieving superior classification accuracy on CIFAR-10/100 and ImageNet while requiring fewer training epochs. Moreover, VarCon exhibits enhanced robustness to hyperparameters, reduced dependency on large batches, stronger low-data performance, and improved corruption resilience, all while maintaining better cross-domain transferability. Beyond empirical advantages, VarCon bridges the theoretical gap between distinguishing and generative paradigms by endowing contrastive objectives with explicit likelihood semantics.
@@ Appendix A Discussion
A.1 Limitation
While Variational Supervised Contrastive Learning (VarCon) demonstrates significant improvements over existing methods, we acknowledge several limitations. First, our approach relies on supervised learning, which limits its applicability where labeled data is scarce or expensive to obtain. Second, although VarCon reduces comparison complexity from quadratic to nearly linear through class-level centroids, computing these centroids still introduces overhead for datasets with numerous classes. Third, cross-domain adaptation may require recalibration of our confidence-adaptive temperature mechanism when facing substantial distribution shifts between source and target domains.
Future work could address these limitations by: 1) extending VarCon to self-supervised settings by leveraging augmentation-based positive pairs with a variational formulation of latent pseudo-labels, which would maintain our probabilistic framework while removing the need for explicit supervision; 2) exploring amortized computation of class centroids through memory banks or representative embedding techniques to reduce computational overhead for large-scale applications; and 3) incorporating domain adaptation techniques such as distribution alignment constraints in the variational objective to improve cross-domain robustness. The core variational formulation of VarCon is flexible enough to accommodate these extensions with relatively modest modifications to the underlying framework.
Despite these limitations, VarCon successfully addresses a fundamental challenge in representation learning by providing a principled probabilistic interpretation of contrastive objectives while simultaneously improving empirical performance across diverse benchmarks. Our approach offers a new theoretical perspective that bridges the gap between generative and distinguishing paradigms, opening promising directions for developing more robust, interpretable, and efficient representation learning methods.
A.2 Broader Impact
@@ B.2 Extended Evaluations
Self-supervised
SimCLR $75.30_{\pm 0.05}$
@@
Supervised
Cross-Entropy $83.47_{\pm 0.06}$
@@
On this more rigorous dataset, VarCon achieves 84.12% Top-1 accuracy, outperforming all baseline methods by a statistically significant margin. The performance gap between VarCon and the strongest self-supervised method (SwAV at 81.56%) is significant, showing a solid 2.56% improvement, suggesting that our probabilistic formulation effectively leverages label information to learn more distinctive representations that better align with human perception of image content.
@@ B.3 Embedding Space Evolution Analysis
The visualization demonstrates clear progression from loosely organized clusters at epoch 50 (KNN classifier: 52.74%) to highly structured, well-separated semantic groups by epoch 200 (79.11%). The most significant improvement occurs between epochs 100-150, where the KNN classifier performance jumps from 59.57% to 71.12%, corresponding to the convergence of our ELBO-derived loss function. During this phase, the KL divergence term effectively aligns the auxiliary posterior with the model’s class posterior while ensuring correct class assignments. By epoch 200, VarCon achieves optimal embedding organization with distinct, compact clusters and clear decision boundaries. We employ KNN classification to directly evaluate the quality of pretrained representations without introducing additional trainable parameters that could mask the intrinsic separability of the learned embeddings. The high KNN classifier performance (79.11%) demonstrates that our confidence-adaptive temperature mechanism successfully provides fine-grained control over intra-class dispersion, creating embedding spaces where semantic similarity corresponds to geometric proximity and enabling effective classification through nearest-neighbor search.
@@
Figure 5: Progressive evolution of VarCon embedding space visualization through t-SNE [SimCLR] during training on ImageNet validation set. Our variational formulation demonstrates systematic improvement in semantic organization, with KNN-classifier accuracy increasing from 52.74% at epoch 50 to 79.11% at epoch 200 as clusters become increasingly well-separated and semantically coherent. The confidence-adaptive temperature mechanism enables fine-grained control over intra-class dispersion, resulting in embedding spaces with clear decision boundaries and hierarchical semantic structure that facilitate effective nearest-neighbor classification without additional parameterized classifiers.
@@ B.4 Hierarchical Clustering Results
Table 6 presents comprehensive clustering quality metrics. VarCon consistently outperforms SupCon across all evaluation criteria, with progressive improvements observed through our ablation. The Adjusted Rand Index (ARI), which measures the similarity between cluster assignments and ground-truth labels while correcting for chance, increases from 0.613 (SupCon) to 0.627 (VarCon with $\epsilon=0$) to 0.634 (VarCon). This progression suggests that the discovered clusters align increasingly closer with true semantic categories. Similarly, the purity metric, which quantifies the fraction of samples in each cluster belonging to the most common class, improves from 0.755 to 0.766 to 0.774, indicating that individual clusters contain progressively fewer mixed-class samples.
The Normalized Mutual Information (NMI) score, measuring the mutual dependence between cluster assignments and true labels, increases from 0.888 to 0.892 to 0.895. This gain reflects a stronger statistical relationship between the discovered structure and semantic categories. The V-measure, which provides the harmonic mean of homogeneity and completeness, shows balanced improvements (0.888 to 0.891 to 0.895). Specifically, homogeneity, measuring whether each cluster contains only members of a single class, improves from 0.886 to 0.890 to 0.893, while completeness, measuring whether all members of a class are assigned to the same cluster, increases from 0.890 to 0.892 to 0.896. These metric improvements align with our theoretical framework. Notably, VarCon with fixed temperature ($\epsilon$) already outperforms SupCon, demonstrating that the ELBO-derived loss itself enhances representation quality. The full VarCon with confidence-adaptive temperature mechanism further improves performance by applying stronger supervision on difficult samples while relaxing constraints on well-classified examples. This adaptive behavior, combined with the KL divergence term that prevents distributional collapse, appears to preserve more nuanced semantic relationships during training. The clustering analysis thus provides empirical support for our variational formulation’s ability to maintain meaningful structure in the representation space, with both components contributing to the superior performance.
@@ Appendix C Theoretical Analysis
$\frac{\partial\mathcal{L}_{\text{VarCon}}}{\partial\epsilon}$ $=[2p_{\theta}(r\mid\bm{z})-1]\times\Bigg\{$
@@
$\quad\times\left\{\sum_{r^{\prime}\neq r}\log\frac{q_{\phi}(r^{\prime}\mid\bm{z})}{p_{\theta}(r^{\prime}\mid\bm{z})}-(C-1)\log\frac{q_{\phi}(r\mid\bm{z})}{p_{\theta}(r\mid\bm{z})}\right\}.$ (46)
@@
  • For confident samples ($p_{\theta}(r\mid\bm{z})>0.5$): The factor $2p_{\theta}(r\mid\bm{z})-1]>0$. When $p_{\theta}(r\mid\bm{z})$ is high (confident prediction), $\tau_{2}$ approaches $\tau_{1}+\epsilon$, making $q_{\phi}$ more uniform. The gradient direction is determined by the aggregate log-ratio term $\sum_{r^{\prime}\neq r}\log\frac{q_{\phi}(r^{\prime}\mid\bm{z})}{p_{\theta}(r^{\prime}\mid\bm{z})}-(C-1)\log\frac{q_{\phi}(r\mid\bm{z})}{p_{\theta}(r\mid\bm{z})}$. This uniform $q_{\phi}$ typically satisfies $q_{\phi}(r\mid\bm{z})<p_{\theta}(r\mid\bm{z})$ for the ground-truth class and $q_{\phi}(r^{\prime}\mid\bm{z})>p_{\theta}(r^{\prime}\mid\bm{z})$ for other classes, making the aggregate term positive. Consequently, the gradient pushes $\epsilon$ to increase, which in turn increases $\tau_{2}$, making $q_{\phi}$ even more uniform and effectively reducing supervision strength for already well-classified samples. This self-reinforcing mechanism prevents overfitting on easy samples while maintaining stable learning dynamics.
  • For difficult samples ($p_{\theta}(r\mid\bm{z})<0.5$): The factor $2p_{\theta}(r\mid\bm{z})-1]<0$. When the model is uncertain, $\tau_{2}$ approaches $\tau_{1}-\epsilon$, creating a sharper distribution. The gradient behavior reverses, causing $\epsilon$ to adjust such that $\tau_{2}$ decreases. This reduction in $\epsilon$ leads to a smaller $\tau_{2}$, creating a sharper $q_{\phi}$ distribution that provides stronger supervision precisely where the model needs it most. The same aggregate log-ratio term that increases $\epsilon$ for confident samples now decreases it due to the negative sign factor. The magnitude of this effect scales with the degree of distributional mismatch, ensuring proportional adaptation to the model’s uncertainty.